OpenStudy (anonymous):

what is the slope between the points (5,3)and (2,-4)

OpenStudy (amistre64):

use the formula ...

OpenStudy (amistre64):

or, subtract one point from the other and stack y/x

OpenStudy (anonymous):

\[slope=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]

OpenStudy (anonymous):

the slope is -7/3 i'm not sure

OpenStudy (anonymous):

\[\frac{ -4-3 }{ 2-5 }=\frac{ -7 }{ -3 }\]

OpenStudy (anonymous):

Do you need help with anything else @dpewu

OpenStudy (anonymous):

He went offline.. :(

OpenStudy (anonymous):

Ridiculous, gets an answer and then leaves -_-

OpenStudy (anonymous):

I deleted my answer because yours explain it better with your fancy \[letters\]

OpenStudy (anonymous):

lol